Question 1: One object is traveling with a constant velocity of 3 m/s. Another is traveling with a constant acceleration of 3
m/s? What is the difference between the motion of these two objects? Explain your answer clearly. Your goal is
to show me how well you understand velocity and acceleration.
1
Expert's answer
2020-10-19T13:22:59-0400

In the first case the distance travelled is 3 meters each second.

In the second case the distance travelled increased by 3 meters each second:


"3\ m,\\3+3=6\ m,\\6+3=9\ m,\\9+3=12\ m..."